OpenStudy (anonymous):

Rewrite in simplest rational exponent form √x • 4√x. Show each step of your process.

OpenStudy (anonymous):

some one please help

OpenStudy (anonymous):

please help @pooja195 and @texaschic101

jimthompson5910 (jim_thompson5910):

is it \[\Large \sqrt{x}*4\sqrt{x}\] OR is it \[\Large \sqrt{x}*\sqrt[4]{x}\]

OpenStudy (anonymous):

the second one

jimthompson5910 (jim_thompson5910):

\[\LARGE \sqrt{x}*\sqrt[4]{x} = \sqrt[2]{x}*\sqrt[4]{x}\] \[\LARGE \sqrt{x}*\sqrt[4]{x} = x^{1/2}*x^{1/4}\] \[\LARGE \sqrt{x}*\sqrt[4]{x} = ???\]

jimthompson5910 (jim_thompson5910):

do you see how to finish up?

OpenStudy (anonymous):

x^1/8

jimthompson5910 (jim_thompson5910):

you don't multiply the exponents, you add them

OpenStudy (anonymous):

x^3/4

jimthompson5910 (jim_thompson5910):

good, the answer is \[\LARGE x^{3/4}\]

OpenStudy (anonymous):

oh thank you so much

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

oh wait one more question

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (anonymous):

Rewrite in simplest radical form x 5 6 x 1 6 . Show each step of your process.

jimthompson5910 (jim_thompson5910):

draw it out please

jimthompson5910 (jim_thompson5910):

or use the equation editor

OpenStudy (anonymous):

\[\frac{ x^5/6 }{ ?x^1/6 }\]

jimthompson5910 (jim_thompson5910):

so \(\LARGE x^{5/6}\) all over \(\LARGE x^{1/6}\) ??

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

the bases are both x so we subtract exponents (top - bottom)